OpenStudy (anonymous):

for f(x) = 4x+1 and g(x) = x^2 - 5 find [g/f](x)

OpenStudy (anonymous):

super confused please help!

OpenStudy (anonymous):

would it just be the two equations over each other?

OpenStudy (jdoe0001):

pretty much, yes

geerky42 (geerky42):

\(\dfrac{g}{f}(x) = \dfrac{g(x)}{f(x)}\)

OpenStudy (anonymous):

so it's x^2-5 / 4x+1 which is D?

OpenStudy (aum):

Make sure the denominator does not get to zero by excluding the x value that would make it zero.

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

nope :(

geerky42 (geerky42):

look closely at A

OpenStudy (anonymous):

yep it was A...

OpenStudy (jdoe0001):

hmmm actaully listen to @aum above... you need to take into account the "constraints"

geerky42 (geerky42):

can x be -1/4? I mean this will get denominator to be 0

OpenStudy (anonymous):

ahh ok so when doing the function problems the denominator can never be zero? Why is that?

OpenStudy (jdoe0001):

well because say 25/0 =?

OpenStudy (jdoe0001):

0/25 = 0 but 25/0 = ?

OpenStudy (anonymous):

undefined. got it. makes sense now! thanks!

OpenStudy (jdoe0001):

\(\bf \cfrac{0}{anything}=0\qquad \cfrac{anything}{0}=undefined\)

OpenStudy (anonymous):

thank you!